# HI-TransPA: Hearing Impairments Translation Personal Assistant

### v1

@@ 5 Conclusion
In this paper, we presented HI‑TransPA, the first instruction‑driven audio‑visual Omni‑Model designed to assist people with hearing loss through natural dialogue and accurate translation. The model integrates speech and high‑frame‑rate lip cues via a redesigned visual encoder and a quality‑guided curriculum learning strategy, enabling robust multimodal understanding under real‑world conditions. Experiments on the purpose‑built HI‑Dialogue dataset demonstrate clear advantages over state‑of‑the‑art ASR systems, LALMs, and general Omni‑Models.
By bridging assistive technology and general‑purpose Omni‑Model research, HI‑TransPA highlights the potential of instruction‑tuned multimodal intelligence for inclusive and accessible communication. In future work, we aim to extend this framework to hearing‑impaired speech training and rehabilitation, contributing to improved accessibility and social inclusion for people with hearing loss.

### v2

@@ 5 Conclusion
We presented HI-TransPA, an instruction-driven audio-visual Omni-Model that designed to assists people with hearing loss in natural dialogue and precise speech translation. With jointly modeling speech and high-frame-rate lip cues with a redesigned visual encoder and a quality-aware curriculum strategy for training, HI-TransPA achieves outstanding performance on multimodal real-world scenarios. Extensive experimental results on the HI-Dialogue dataset demonstrate that HI-TransPA consistently outperforms existing ASR methods, LALMs, and general Omni-Models, highlighting the potential of instruction-tuned multimodal intelligence for inclusive communication.